CNNBC recently reported that the mean annual cost of auto insurance is 1038 dollars. Assume the standard deviation is 238 dollars. You take a simple random sample of 74 auto insurance policies. Find the probability that a single randomly selected value is less than 991 dollars. P(X < 991) = Find the probability that a sample of size n = 74 n = 74 is randomly selected with a mean less than 991 dollars. P(M < 991) = Enter your answers as numbers accurate to 4 decimal places.
Answer #1

Mean annual cost of auto insurance is 1038 dollars. That is \mu = 1038

Given, standard deviation is 238 dollars. So, \sigma = 238

Assuming the annual cost of auto insurance follows Normal distribution,

Probability that a single randomly selected value is less than 991 dollars = P(X < 991)

= P[Z < (991 - 1038)/238]

= P[Z < -0.1975]

= 0.4217 (Using Z table)

Standard error of mean = \sigma / \sqrt{n} = 238/ \sqrt{74} = 27.6669

Probability that mean less than 991 dollars. P(M < 991)

= P[Z < (991 - 1038)/27.6669]

= P[Z < -1.6988]

= 0.0447 (Using Z table)